Grade 12Algebra

Solve:
4 + (7/8)
Solve:
4 + (7/8)
Solve:
4 + (7/8)
Solve:
4 + (7/8)

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

For addition of two unlike fractions, first change them to the like fractions.
LCM of 1, 8 = 8
Now, let us change each of the given fraction into an equivalent fraction having 8 as the denominator.
= [(4/1) × (8/8)] = (32/8)
= [(7/8) × (1/1)] = (7/8)
Now,
= (32/8) + (7/8)
= [(32 + 7)/8]
= (39/8)
= 4 whole 7/8

Thanks